OpenStudy (anonymous):

sam has scores of 96 and 86 on his first two geometry tests. What possible scores can he make on his third test so that is average is at least 90?

OpenStudy (anonymous):

For my final answer I get x>=88 is this correct?

Parth (parthkohli):

The average is the mean here. Let's say that it's supposed to be x to make it at least 90. Inequality here. \(\Large \color{Orange}{\Rightarrow {96 + 86 + x \over 3} \ge 90 }\) \(\Large \color{Orange}{\Rightarrow {182 + x} \ge 270 }\) \(\Large \color{Orange}{\Rightarrow x \ge 88 }\) Correct :)

OpenStudy (anonymous):

Thanks

OpenStudy (compassionate):

\[x \ge 90\]

Parth (parthkohli):

@Compassionate no! NO!

Parth (parthkohli):

Just see my answer.

OpenStudy (compassionate):

Sorry, Parth. You're correct.